how to determine order of a reaction?tell me in detail.

Grade:12

2 Answers

deepak kumar singh AskiitiansExpert-IITM
13 years ago

dear shamshad reyaz

first of all remember one thing about the order of reaction: order of reaction is a practically measured value, one can not determine it just by the reaction's equation, but still thr ways to determine the order of reaction.

In chemical kinetics, the order of reaction with respect to a certain reactant, is defined as thepower to which its concentration term in the rate law or rate equation is raised.

For example, given a chemical reaction 2A + B → C with a rate equation : r = k[A]α[B]β

the reaction order with respect to A would be α and with respect to B would be β, the total reaction order would be α+β. It is not necessary that the order of a reaction be a whole number – zero and fractional values of order are possible. Reaction orders can be determined only by experiment. Their knowledge allows conclusions about the reaction mechanism.

The reaction order is not necessarily related to the stoichiometry of the reaction, unless the reaction is elementary. Complex reactions may or may not have reaction orders equal to their stoichiometric coefficients.

----->supoose if above reaction would be elementay then the order with repect to A and B would be 2 and 1 repectively and overall ordr of reaction would be 2+1=3.

------>suppose if a question provide you some experimental data about the concentration of reactant , then you have to put those values in the rate equation and find out the valuse of α and β.

-------> another approach to calculate the order of reaction by the unit of rate constant "k".

          unit of rate constant (k) = mol1-n·Ln-1·s-1 

where n is the order of reaction.

 but as i think the best way to calculate the order of reaction is to first of all understand the concept behind the order of reaction after that just appy that concept and you will youself find out the way to determine the order.

after all of this there are so many methods to determine the value of order of reaction.

so first of all read the concept and after that if you have some problem you can ask us.

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum..

ALL THE BEST

regard

Deepak kumar singh

B.TECH IIT madras

Srinivas Rao
13 years ago

Experimentally order of a reaction can be detrmined by any one of the following four methods:

Vant Hoff differential method: By measuring the rate of reaction at two different concentrations; the order of reaction n can be calculated:

n = (log rate-1 at conc C1 - log rate-2 at conc C2) / (log C1 - log C2)

Half-time method: The half-life period of a reaction t1/2 proportional to 1/a(n-1) k where ‘n’ is the order of reaction. When the initial concentration is a1, the half-life period be 1t1/2 When the initial concentration is a2, the half-life period be 2t1/2 Then, 1t1/2 / 2t1/2 = (a2 /a1) (n-1) By knowing 1t1/2 , 2t1/2 , a1 and a2 the value of ‘n’, the order of reaction is calculated. Ostwald’s isolation method: The order of reaction w.r.t. a reactant is calculated by comparing the rates of the reaction at two different concentrations; keeping the concentration of others a constant. This process is repeated for all the reactants to find out the over all order of reaction. For example, Rate = [A]x [B]y -----(1) and 4 Rate = [2A]x [B]y -----(2) Then, dividing equation (2) with (1) , we get 4 = 2x or x = 2, Second order Similarly, if Rate = [A]x [2B]y -----(3) then, dividing equation (3) with (1) we get 1 = 2y or y = 0, Zero order. So, the total order of the reaction is 2+0=2.

Trial and error method or integrated rate equation method: At various time intervals, the corresponding concentration of reactants is measured. On substituting these values in the integrated rate expressions, a constant value is obtained only for a particular order reaction, and not for any others. Thus, order of a reaction is known by trial and error method. For zero order, k= x/t For first order, k= (2.303/t) log (a / (a-x))